OpenStudy (anonymous):

differentiate f(x) = (lnx)/(1-x)

OpenStudy (cwrw238):

use the quotient rule

OpenStudy (cwrw238):

d(u/v)/dx = v*du/dx - u * dv/dx ------------------- v^2

OpenStudy (anonymous):

i used product rule and i got (x-x^2+lnx)/(1-x^2) i just want to check

OpenStudy (anonymous):

i let u=lnx v=(1-x)^-1 du/dx=1/x and dv/dx=1/(1-x)^2

OpenStudy (cwrw238):

oh ok that's lnx (1-x)^-1 derivative is lnx * -(1-x)-2 *-1 + (1-x)^-1 * (1/x)

OpenStudy (anonymous):

ya,so im kinda unsure coz im suppose to find f'(1). but i get 0/0

OpenStudy (anonymous):

\[f'(x) =\frac{ x-x^{2}+lnx }{ (1-x)^{2} }\]

OpenStudy (cwrw238):

hmm i'll try simplifying...

OpenStudy (cwrw238):

I get x lnx - x + 1 ----------- x(1 - x)^2

OpenStudy (anonymous):

hm,, it will still get 0/0 so the answer for f'(1) =0?

OpenStudy (cwrw238):

yes still 0/0 - no 0/0 in indeterminate

OpenStudy (cwrw238):

i'll just recheck my working...

OpenStudy (anonymous):

i see, okay thanks!

OpenStudy (cwrw238):

yes I still get same value for f'(x) using the quotient rule so f'(1) is indeterminate

OpenStudy (anonymous):

okay thank you very much!(:

OpenStudy (cwrw238):

yw as it happened it was easier to use the quotient rule in this case